A pool can be filled by one pipe in 4 hours and by a second pipe in 12 hours.how long will it take using both pipes to fill the pool?
Answer by richwmiller(17219)

1/4+1/12=1/x
3/12+1/12=1/x
4/12=1/x
1/3=1/x
x=3